OpenStudy (anonymous):

Line m has slope -2 and contains the point (3, -7). Which of the following is a point-slope equation of line m?

OpenStudy (saifoo.khan):

\[y - y_1 = m(x-x_1)\]

OpenStudy (saifoo.khan):

\[y - (-7) = -2 (x - 3)\]Solve until you get it in; y = mx + b form

OpenStudy (anonymous):

soo how would i have too do it

OpenStudy (saifoo.khan):

solve for "y".

OpenStudy (anonymous):

i really dk how to do this

OpenStudy (saifoo.khan):

@hoblos , please help here. i will brb.

OpenStudy (hoblos):

the general point-slope equation of line of slope m and passing by (x1,y1) is y-y1=m(x-x1) here m=-2 x1=3 y1=-7 so y-(-7) = -2(x-3) y+7=-2(x-3)